A ball is projected from a certain point on the surface of planet at an angle with horizontal surface. the horizontal and vertical displacement x and y vary with time in second as : x=10√3t and y=10t-t^2.the maximum height attained isA.100m. B.75m. C.50m. D.25m

Grade:11

6 Answers

Kshitij Sharma
7 years ago
D.25m is the correct answer.Using vector equation \vec{r}=\vec{u}t+\frac{1}{2}\vec{{g}`}t^{2}​ where {g}`​ is acc. due to grav. on the planet, we get :  \vec{r}=u\cos \theta t\hat{i}+(u\sin \theta t-{g}`t^{2})\hat{j}​ = 10\sqrt{3}t\hat{i}+(10t-t^{2})\hat{j}​ . On comparing and evaluating corresponding terms, we get \tan \theta =\frac{1}{\sqrt{3}}   so,  \theta =30^{\circ}​ , |u|=20m/s and |{g}`|=2ms^{-2} .From y_{max}=\frac{u^{2}\sin ^{2}\theta }{2{g}`}​, max. height=25 m.     
utpal raj
7 years ago
From  x=10√3t and y=10t-t^2
substitute the t value in y eqn. we get,
y = x/root(3) – x^2/300
on comparing with the trajectory eqn. of motion we get 
thetha = 30 degree and on solving for we get u = 20root(5) m/s
Maimum height = u^2sin^2thetha/g => 2000/4g => 50 meter.
Gaurav
6 years ago
From x=10√3t and y=10t-t²substitute the t value in y eqn. we get,y = x/√3– x²/300on comparing with the trajectory eqn. of motion we get thetha = 30 degree and on solving for we get u²=2000Height= u²sin²30/2gH=2000/8gH=25
Animesh Agarwal
5 years ago
y = 10t - t2vy = dy/dt = 10- 2tAt maximum height, vy = 0m/sTime taken to reach the maximum height, t = 5 secondsMaximum height, y = 10(5) - (5)2 = 25 m
Jayakeerti
5 years ago
By differentiating y = 10t - t^2 we get vertical component of velocity V = 10 -2t. 
At max height. V = 0, thus we get t = 5s.
By substituting it at y = 10t -t^2, we get y = 25m
Kharanshu arora
3 years ago
x=ucos€.t.                        Y=usin€.t-g/2t²
x=10√3t.   ----(given).      Y=10 t-t².  -----(given)
 
Thus,
ucos£=10√3.          usin£=10.     g/2=1
                                                    g=2
 
 
Hmax= (usin£)²/2g
          =(10)²/2(2)
          =100/4
          = 25 m